OpenStudy (anonymous):

solve 13-y is less than or equal to 2(y-4)

OpenStudy (goldphenoix):

Does it look like: \[\large \large 13-y \le 2(y-4)\]?

OpenStudy (anonymous):

yes expcpt the inequality should be greater than sorry miss typed it greater than or equal to

OpenStudy (goldphenoix):

Oh okay.\[\large \large 13-y \ge2(y-4)\] The first step is to do the distributive property on 2(y-4) So, 2 * y = ? 2 * -4 = ?

OpenStudy (anonymous):

2y -6

OpenStudy (goldphenoix):

2 * y does equal to 2y. But what does 2 * -4 give you?

OpenStudy (goldphenoix):

If you are not sure, then what does 2 multiply by 4 give you?

OpenStudy (anonymous):

-8

OpenStudy (goldphenoix):

Great. 2 * -4 = -8. So now your equation should look like: \[\large \large 13 - y \ge 2y - 8\] Next step is to move the variable to one side of the equation, and the number to the other side of the equation. So, we can add y to both side. -y + y = ? 2y + y = ?

OpenStudy (anonymous):

3y

OpenStudy (goldphenoix):

Great! 2y + y = 3y. I guess you already know that -y + y cancel out. So now you have: \[\large \large 13 \ge 3y - 8\] Do you know what to do next?

OpenStudy (anonymous):

add eight which gives 21 divded by 3 equals 7

OpenStudy (goldphenoix):

Good job! So, \[\large \large y \ge 7\]

OpenStudy (anonymous):

thank you so much im having trouble with another could you pleaseee helpme